Bottle A, containing 12L of 15% acid, is combined with bottle B, containing 3L of 25% acid. Bottle C is 26% acid. How much of each solution is neede to have 24L of a 20% acid solution?
Answer by ankor@dixie-net.com(22740)

Bottle A, containing 12L of 15% acid, is combined with bottle B, containing 3L of 25% acid.
Bottle C is 26% acid.
How much of each solution is needed to have 24L of a 20% acid solution?
:
Let C = amt of 26% acid solution required
:
.26C + .15(12) + .25(3) = .20(24)
Do the math
.26C + 1.8 + .75 = 4.8
.26C + 2.55 = 4.8
.26C = 4.8 - 2.55
.26C = 2.25
C = 2.25/.26
C = 8.654 L of 26% acid
:
"How much of each solution is needed to have 24L of a 20% acid solution?"
12L of A
3 L of B
8.654L of C